# MasterWeaver: Taming Editability and Identity for Personalized Text-to-Image Generation

### v2

@@ Abstract
Text-to-image (T2I) diffusion models have shown significant success in personalized text-to-image generation, which aims to generate novel images with human identities indicated by the reference images. Despite promising identity fidelity has been achieved by several tuning-free methods, they usually suffer from overfitting issues. The learned identity tends to entangle with irrelevant information, resulting in unsatisfied text controllability, especially on faces. In this work, we present MasterWeaver, a test-time tuning-free method designed to generate personalized images with both faithful identity fidelity and flexible editability. Specifically, MasterWeaver adopts an encoder to extract identity features and steers the image generation through additional introduced cross attention. To improve editability while maintaining identity fidelity, we propose an editing direction loss for training, which aligns the editing directions of our MasterWeaver with those of the original T2I model. Additionally, a face-augmented dataset is constructed to facilitate disentangled identity learning, and further improve the editability. Extensive experiments demonstrate that our MasterWeaver can not only generate personalized images with faithful identity, but also exhibit superiority in text controllability. Our code will be publicly available at https://github.com/csyxwei/MasterWeaver.
@@
Figure 1: With one single reference image, our MasterWeaver can generate photo-realistic personalized images with diverse clothing, accessories, facial attributes and actions in various contexts. In comparison with existing methods, our method exhibits superior editability while maintaining faithful identity fidelity.
@@ 1 Introduction
Recently, text-to-image diffusion models [65] have demonstrated superior capabilities in generating high-quality and creative images. Building upon these advancements, personalization methods [20, 66, 82, 87] further enable the generation of a specific visual subject (e.g., objects, animals, or people) as indicated by several reference images. In this work, we focus on the personalization of human identities, which has wide applications, such as personalized portrait photos [40], art creation [75], and visual try-on [77].
@@
Earlier studies, e.g., Dreambooth [66] and Textual Inversion [20], usually learned the novel identities from reference images by optimizing the word embeddings or model parameters. Albeit the promising results, these methods require per-identity optimization, which is time-consuming and impractical for real applications. Additionally, they often take multiple images for identity learning and suffer from poor editability in limited-data scenarios (see the top of Fig. 1). Recent tuning-free methods, e.g., FastComposer [84] and IP Adapter [87] trained additional visual encoders to extract the identity information, and inject them through model tuning or adapters. After training on human datasets, these methods could produce personalized results in several denoising steps with only one identity image as the reference. Despite these methods having achieved promising identity fidelity, they usually suffer from overfitting issues. Since the reference identity image and the input image used during training are from the same image, the methods tend to directly copy the reference image during generation, resulting in poor editability (see the top of Fig. 1). Photomaker [40] took multiple images of the same identity to learn a stacked face embedding, which can be used to generate images with diverse attributes. However, its identity fidelity is heavily sacrificed for improved editability.
@@
To address the above issue, we propose MasterWeaver, a tuning-free method for generating personalized images with both faithful identity fidelity and flexible text controllability. Following [87, 82], MasterWeaver adopts an encoder to extract the identity features and steers the image generation through additional introduced cross attention. To improve the editability while keeping the identity fidelity, we propose identity-preserved editability learning. We first propose an editing direction loss to facilitate the model training. Specifically, we identify the editing direction in the feature space of diffusion model by inputting paired text prompts that denote an editing operation, e.g., (a photo of a person, a photo of a person with <attribute>). As the direction’s computation solely depends on the attribute difference, such a direction captures the meaningful semantic editing prior and is unrelated to identity. By aligning the editing direction of our MasterWeaver with that of original T2I model, we can significantly enhance our method’s text controllability without compromising identity fidelity.
@@
Additionally, we construct a face-augmented dataset to facilitate disentangled identity learning. In particular, we utilize the face editing method [50] to modify the attributes of reference identity images, which are then incorporated into our augmented dataset. In this dataset, the reference identity image and corresponding input image have the same identity but differ in a single attribute (e.g., hair color, style, or expression). Such a controlled attribute misalignment can effectively facilitate model learning to extract faithful identity features disentangled with attribute details, thereby improving editability. As illustrated in Fig. 1, with only one reference image, our MasterWeaver can generate photo-realistic personalized images with faithful identity and diverse contexts.
@@
Extensive experimental evaluations demonstrate that our MasterWeaver outperforms current state-of-the-art methods. The contributions of this work are summarized as follows:
@@
  • We propose the MasterWeaver, a novel method that can generate personalized images with high efficiency, faithful identity fidelity, and flexible controllability.
@@
  • Experimental results show that our MasterWeaver can generate the target identity faithfully, while showing more flexible text controllability.
@@ 2.1 Text-to-Image Diffusion Models
Recently, diffusion models [26, 16] have demonstrated remarkable capabilities in generating photo-realistic images and have been widely adopted in text-to-image (T2I) generation [5, 69, 63, 17, 53, 65, 7, 57]. Benefiting from the advance in language model [71, 62, 61] and extensive text-image datasets [71], these text-to-image diffusion models are capable of generating textually coherent and high-quality images. Among them, Stable Diffusion [65] is one of the representative open-sourced models, which performs a diffusion process in the latent space to reduce computational complexity. It has demonstrated the superior capacity in generating high-quality and diverse images, and facilitates a surge of recent advance in downstream tasks [89, 92, 58, 49]. Stable Diffusion XL (SDXL) [57] employed a larger UNet and an additional text encoder to achieve superior image generation quality and enhanced textual fidelity. We also employ the Stable Diffusion as the text-to-image model in our experiments.
@@ 2.2 Personalized Text-to-image Generation
Building upon the advances in text-to-image models, personalization methods [20, 66, 82, 87] further enable the generation of specific visual concepts (e.g., objects, animals, or people) as indicated by several reference images. Earlier studies [20, 18, 66, 34, 81, 19, 96, 29, 52, 10, 36, 8, 60, 45, 46, 90, 27, 94, 4, 76, 55, 74, 68, 48, 3, 83, 25] usually learned the novel concept from reference images by optimizing word embeddings or model parameters. For example, Textual Inversion [20] optimized a new “word” embedding using a few reference images to learn the target concept. DreamBooth [66] finetuned the parameters of the T2I model to align the unique identifier with the high-fidelity new concept. To improve the computation efficiency, Custom Diffusion [34] proposed only updating the key and value mapping parameters in cross attention layers. CelebBasis [88] projected the human identity into a celeb space, extracted from celebrity identities, and shown superior editability. Although the results are promising, these methods usually take multiple images to learn the concept and require substantial time for finetuning. To reduce the tuning time, several methods [21, 22] suggested initial pre-training on large datasets followed by tuning on smaller datasets to expedite the process. However, it still takes tens of steps for finetuning, limiting its practicality in real applications.
Recent tuning-free methods [73, 2, 59, 37, 82, 93, 51, 12, 31, 9, 28] train additional visual encoders to extract the concept information, and inject them through model tuning or adapters. By training on personalization datasets, these methods could produce personalized results in tens of denoising steps with only one reference image. For example, ELITE [82] employed a global mapping encode the CLIP [61] features of concept image as textual embedding, and further improves finer details with a local mapping network. BLIP Diffusion utilized a pre-trained BLIP2 [38] as the feature extractor to produce a visual representation aligned with the text. Additionally, several methods [79, 67, 85, 75, 86, 41, 30, 11, 56, 78, 44, 39, 13] are designed for personalization of human identities. DreamIdentity [13] crafted an editing dataset based on celebrity identity and a pre-trained stable diffusion model for editability learning. Photomaker [40] collected a human dataset that consists of multiple images of the same identity and used them to learn a stacked face embedding, which could generate images with diverse attributes.
In this work, we propose the tuning-free MasterWeaver. Compared with existing methods, our MasterWeaver could generate personalized images with high efficiency, faithful identity preservation, and flexible controllability.
@@ 2.3 Face Editing
Face generation and editing are always popular in computer vision and computer graphics. With the development of Generative Adversarial Networks (GANs) [24], several methods have been proposed for high-quality face generation [32, 33] and the flexible face editing [50, 72, 6, 42, 54]. Generally, existing GAN-based face editing methods can be classified into two types. The first type of methods [14, 43, 15] utilized image-to-image translation techniques, where the original face image is also fed to the network to produce the edited images. The second type of methods [72, 6, 42] first inverted the given face into the latent code of pre-trained GANs [80, 1, 64], and then manipulated it with specific directions. Recently, with the advance of CLIP [61], several methods [50, 54, 23, 91] are proposed to perform the face editing based on text prompts.
In our work, we adopt the face editing method [50] to construct a face-augmented dataset, which can be used for disentangled identity training.
Figure 3: Illustration of Editing Direction Loss. By inputting paired text prompts that denote an editing operation, e.g., (a photo of a woman, a photo of a smiling woman), we identify the editing direction in the feature space of diffusion model. Then we align the editing direction of our MasterWeaver with that of original T2I model to improve the text controllability without affecting the identity.
@@ 3 Proposed Method
Given one single reference image $x$ of a specific identity, personalized text-to-image generation aims to generate photo-realistic images of the given identity according to the text prompts $y$. Nonetheless, the generated images are expected to keep a faithful identity and exhibit diversity in attributes, actions, contexts, and so on. To achieve this goal, we propose MasterWeaver, a tuning-free method for generating personalized images with promising identity fidelity and flexible editability. As illustrated in Fig. 2, our MasterWeaver first employs an identity mapping network to encode identity feature, and incorporates it with text to guide the image generation. To improve the model’s editability while keeping identity fidelity, we further propose the id-preserved editability learning, including an editing direction loss and a face-augmented dataset. In the following, we first present an overview of the text-to-image model utilized in our approach (Sec. 3.1). Then, we will introduce the details of the face identity injection (Sec. 3.2) and id-preserved editability learning (Sec. 3.3). Finally, we will give our learning objective (Sec. 3.4).
@@ 3.1 Preliminary
In this work, we employ the large-scale pre-trained Stable Diffusion [65] as our text-to-image model, which can generate diverse and photo-realistic images. It consists of two components, i.e., the autoencoder $\mathcal{E}(\cdot)$, $\mathcal{D}(\cdot)$) and the conditional diffusion model $\epsilon_{\theta}(\cdot)$. Specifically, the encoder $\mathcal{E}(\cdot)$ is trained to map an image $x$ to a lower dimensional latent space $z=\mathcal{E}(x)$, and the decoder $\mathcal{D}(\cdot)$ is trained to map the latent code back to the image so that $D(\mathcal{E}(x))\approx x$. The conditional diffusion model $\epsilon_{\theta}(\cdot)$ is trained on the latent space of autoencoder, which can generate latent codes based on text condition $y$. The mean-squared loss is employed to train the model:
@@
Cross attention is adopted in Stable Diffusion to steer the generation process by text prompt. Specifically, cross attention adopts the latent image feature $h$ and text feature $\tau(y)$ as input, and transforms them to query $Q=W_{Q}\cdot h$, key $K=W_{K}\cdot\tau(y)$ and value $V=W_{V}\cdot\tau(y)$ by projection layers. $W_{Q}$, $W_{K}$, and $W_{V}$ are weight parameters of query, key, and value projection layers, respectively. Then, attention maps are computed with:
@@ 3.2 Faithful Identity Injection
To leverage the pre-trained stable diffusion for generating the personalized images of specified identity, we first introduce a mapping network to encode the identity features, and then integrate the encoded features into the image generation process. As shown in Fig. 2 (b), we employ the pre-trained CLIP image encoder $\psi(\cdot)$ to extract the identity information. To ensure identity fidelity, we utilize the local patch image feature from the penultimate layer of the CLIP model, which contains rich details and is sufficient to represent the target identity faithfully. To bridge the gap between the CLIP features and the stable diffusion model, we further employ an identity (ID) mapper $M(\cdot)$, which is trained to project the CLIP features as identity feature $f$:
@@
To effectively integrate the identity feature $f$ into the generation, we adopt the dual cross attention mechanism [87, 82]: $\text{Attention}(Q,K^{f},V^{f})$, where $K^{f}=W^{f}_{K}\cdot f$ and $V^{f}=W^{f}_{V}\cdot f$ represent the projected key and value matrices of the identity information, respectively. $W^{f}_{K}$ and $W^{f}_{V}$ are two additional introduced projection layers in each cross attention block of stable diffusion, which are optimized with our ID mapper simultaneously. To fuse the identity information with the text information, we sum them by:
@@ 3.3 ID-Preserved Editability Learning
Although the proposed method in Sec. 3.2 can generate personalized images with faithful identity, its text controllability is limited. Since the reference identity image and input image used during training are from the same image (as shown in Fig. 4 left), when we train the model under the reconstruction paradigm, the learned identity feature is inevitably entangled with facial attributes (e.g., hair, pose and expression). Such an entanglement weakens the control of the text over generated images, and the model tends to directly copy reference images during generation, resulting in poor editability and diversity. To improve editability while keeping a faithful identity, we propose id-preserved editability learning, which consists of an editing direction loss and a face-augmented dataset.
Editing Direction Loss. To improve the text controllability, we first propose an editing direction loss to facilitate the model learning. The stable diffusion model is well-recognized for its superior text controllability and can generate images that closely align with provided textual descriptions. Intuitively, we can employ its editing capability to regularize our MasterWeaver. As shown in Fig. 3, by leveraging paired text prompts that denote an editing operation e.g., ($y=$ a photo of a person, $y^{\prime}=$ a photo of a person with <attribute>), we can identify an editing direction in the feature space of diffusion unet model,
@@
Such an editing direction captures the meaningful prior of stable diffusion for semantic editing. Then, we align the editing directions of our MasterWeaver with those of stable diffusion during training,
@@
where $\epsilon_{\theta^{\prime}}$ denotes our MasterWeaver model and $\text{Sim}(\cdot,\cdot)$ denotes the calculation of cosine similarity. $m$ is the mask of the facial region to ensure the loss is only calculated on the facial area. Obviously, if our MasterWeaver neglects the text prompts during generation, its editing directions will be different from those of Stable Diffusion, resulting in a large loss. Besides, the calculation of editing direction is sorely based on the text difference, and the encoded semantic information is unrelated to the identity. Therefore, training with our proposed direction loss can improve the model’s editability significantly, with the moderate sacrifice of the identity fidelity. In our experiments, we have collected several attribute-related prompt pairs for training, including hair, age, expression, etc. More details can be found in the Suppl.
Face-Augmented Dataset Construction. As we analyzed, one of the reasons behind the poor editability is that the learned model entangles the identity with irrelevant attributes and tends to copy the reference image directly, neglecting the text prompt. To address this, we further build a face-augmented dataset designed for disentangled identity training. As illustrated in Fig. 4, we employ the E4E [64] and DeltaEdit [50] to perform the attribute editing for the reference identity image. Then, we construct our face-augmented dataset by combining the input image, text prompt, and edited face images. In this dataset, the reference identity image and corresponding input image have the same identity but differ in one specific attribute (e.g., hair color, style, or expression). Such a controlled attribute misalignment can effectively facilitate model learning to extract faithful identity features disentangled with attribute details, thereby improving editability. To maintain identity fidelity, we filter to exclude edited face images with lower face similarity to reference images. In our experiments, we utilize a mix of face-augmented and original datasets for balanced editability and identity fidelity. More details of the dataset are provided in the Suppl.

### v3

# MasterWeaver: Taming Editability and Face Identity for Personalized Text-to-Image Generation

@@ Abstract
Text-to-image (T2I) diffusion models have shown significant success in personalized text-to-image generation, which aims to generate novel images with human identities indicated by the reference images. Despite promising identity fidelity has been achieved by several tuning-free methods, they often suffer from overfitting issues. The learned identity tends to entangle with irrelevant information, resulting in unsatisfied text controllability, especially on faces. In this work, we present MasterWeaver, a test-time tuning-free method designed to generate personalized images with both high identity fidelity and flexible editability. Specifically, MasterWeaver adopts an encoder to extract identity features and steers the image generation through additionally introduced cross attention. To improve editability while maintaining identity fidelity, we propose an editing direction loss for training, which aligns the editing directions of our MasterWeaver with those of the original T2I model. Additionally, a face-augmented dataset is constructed to facilitate disentangled identity learning, and further improve the editability. Extensive experiments demonstrate that our MasterWeaver can not only generate personalized images with faithful identity, but also exhibit superiority in text controllability. Our code can be found at https://github.com/csyxwei/MasterWeaver.
@@
Figure 1: With one single reference image, our MasterWeaver can generate photo-realistic personalized images with diverse clothing, accessories, facial attributes and actions in various contexts. In comparison with existing methods, our method exhibits superior editability while maintaining high identity fidelity.
@@ 1 Introduction
Recently, text-to-image diffusion models [68] have demonstrated superior capabilities in generating high-quality and creative images. Building upon these advancements, personalization methods [20, 69, 85, 90] further enable the generation of a specific visual subject (e.g., objects, animals, or people) as indicated by one or several reference images. In this work, we focus on the personalized image generation of human identities, which has wide applications, such as personalized portrait photos [42], art creation [78], and visual try-on [80].
@@
Earlier studies, e.g., Dreambooth [69] and Textual Inversion [20], usually learned the novel identities from reference images by optimizing the word embeddings or model parameters. Albeit the promising results, these methods require per-identity optimization, which is time-consuming and impractical for real applications. Additionally, they often take multiple images for identity learning and suffer from poor editability in limited-data scenarios (see the top of Fig. 1). Recent tuning-free methods, e.g., FastComposer [87] and IP Adapter [90] trained additional visual encoders to extract the identity information, and inject them through model tuning or adapters. After training on human datasets, these methods could produce personalized results in several denoising steps with only one identity image as the reference. Despite these methods having achieved promising identity fidelity, they usually suffer from overfitting issues. Since the reference identity image and the input image used during training are from the same image, these methods tend to directly copy the reference image during generation, resulting in poor editability (see the top of Fig. 1). Photomaker [42] took multiple images of the same identity to learn a stacked face embedding, which can be used to generate images with diverse attributes. However, its identity fidelity is sacrificed for improving editability.
@@
To address the above issues, we propose MasterWeaver, a tuning-free method for generating personalized images with both high identity fidelity and flexible text controllability. Following [90, 85], MasterWeaver adopts an encoder to extract the identity features and steers the image generation through additionally introduced cross attention. To improve the editability while keeping the identity fidelity, we propose identity-preserved editability learning. We first propose an editing direction loss to facilitate the model training. Specifically, we identify the editing direction in the feature space of diffusion model by inputting paired text prompts that denote an editing operation, e.g., (a photo of a person, a photo of a person with <attribute>). As the direction’s computation solely depends on the attribute difference, such a direction captures the meaningful semantic editing prior and is unrelated to identity. By aligning the editing direction of our MasterWeaver with that of the original T2I model, we can significantly enhance the text controllability without compromising identity fidelity.
@@
Additionally, we construct a face-augmented dataset to facilitate disentangled identity learning. In particular, we utilize the face editing method [53] to modify the attributes of reference identity images, which are then incorporated into our augmented dataset. In this dataset, the reference identity image and its corresponding input image have the same identity but differ in a single attribute (e.g., hair color, style, or expression). Such a controlled attribute misalignment can effectively facilitate model learning to extract faithful identity features disentangled with attribute details, thereby improving editability. As illustrated in Fig. 1, with only one reference image, our MasterWeaver can generate photo-realistic personalized images with faithful identity and diverse contexts.
@@
Extensive experimental evaluations demonstrate that our MasterWeaver outperforms current state-of-the-art methods. The main contributions of this work are summarized as follows:
@@
  • We propose MasterWeaver, a novel method that can generate personalized images with high efficiency, faithful identity, and flexible controllability.
@@
  • Experimental results show that MasterWeaver can generate the target identity faithfully, while showing more flexible text controllability.
@@ 2.1 Text-to-Image Diffusion Models
Recently, diffusion models [27, 16] have demonstrated remarkable capabilities in generating photo-realistic images and have been widely adopted in text-to-image (T2I) generation [5, 72, 66, 17, 56, 68, 7, 60]. Benefiting from the advance in language model [74, 65, 64] and large-scale text-image datasets [74], these T2I diffusion models are capable of generating textually coherent and high-quality images. Among them, Stable Diffusion [68] is one of the representative open-sourced models, which performs a diffusion process in the latent space to reduce computational complexity. It has demonstrated the superior capacity in generating high-quality and diverse images, and facilitated a surge of recent advances in downstream tasks [92, 94, 61, 44, 31, 52, 95]. Stable Diffusion XL (SDXL) [60] employed a larger UNet and an additional text encoder to achieve superior image generation quality and enhanced textual fidelity. Stable Diffusion is also employed as the T2I model in our experiments.
@@ 2.2 Personalized Text-to-image Generation
Building upon the advances in T2I models, personalization methods [20, 69, 85, 90] further enable the generation of specific visual concepts (e.g., objects, animals, people) as indicated by one or several reference images. Earlier studies [20, 18, 69, 36, 84, 19, 99, 30, 55, 10, 38, 8, 63, 48, 49, 93, 28, 97, 4, 79, 58, 77, 71, 51, 3, 86, 25] usually learned the novel concept from reference images by optimizing word embeddings or model parameters. For example, Textual Inversion [20] optimized a new “word” embedding using a few reference images to learn the target concept. DreamBooth [69] finetuned the parameters of the T2I model to align the unique identifier with the high-fidelity new concept. To improve the computation efficiency, Custom Diffusion [36] only updated the key and value mapping parameters in cross attention layers. CelebBasis [91] projected the human identity into a celeb space, extracted from celebrity identities, and showed superior editability. Though the results are promising, these methods usually take multiple images to learn the concept and require substantial time for finetuning. To reduce the tuning time, several methods [21, 22] suggested initial pre-training on large datasets followed by tuning on smaller datasets to expedite the process. However, they still take tens of steps for finetuning, limiting their practicality.
Recent tuning-free methods [76, 2, 62, 39, 85, 96, 54, 12, 33, 9, 29] train additional visual encoders to extract the concept information, and inject them via model tuning or adapters. By training on personalization datasets, these methods could produce personalized results in tens of denoising steps with only one reference image. For example, ELITE [85] employed a global mapping to encode the CLIP [64] features of concept image as textual embedding, and further improved finer details with a local mapping network. BLIP Diffusion utilized a pre-trained BLIP2 [40] as the feature extractor to produce a visual representation aligned with the text. Additionally, several methods [82, 70, 88, 78, 89, 43, 32, 11, 59, 81, 47, 41, 13] are designed for personalization of human identities. DreamIdentity [13] crafted an editing dataset based on celebrity identity and a pre-trained stable diffusion model for editability learning. Photomaker [42] collected a human dataset that consists of multiple images of the same identity and used them to learn a stacked face embedding, which could generate images with diverse attributes.
In this work, we propose the tuning-free MasterWeaver. Compared with existing methods, MasterWeaver could generate personalized images with high efficiency, faithful identity preservation, and flexible controllability.
@@ 2.3 Face Editing
Face generation and editing are popular topics in computer vision and computer graphics. With the development of Generative Adversarial Networks (GANs) [24], several methods have been proposed for high-quality face generation [34, 35] and flexible face editing [53, 75, 6, 45, 57]. Generally speaking, existing GAN-based face editing methods can be classified into two types. The first type of methods [14, 46, 15] utilize image-to-image translation techniques, where the original face image is fed to the network to produce the edited image. The second type of methods [75, 6, 45] first invert the given face into the latent code of pre-trained GANs [83, 1, 67], and then manipulate it with specific directions. Recently, with the advance of CLIP [64], several methods [53, 57, 23] have been proposed to perform face editing based on text prompts.
In our work, we adopt the face editing method in [53] to construct a face-augmented dataset, which can be used for disentangled identity training.
Figure 3: Illustration of Editing Direction Loss. By inputting paired text prompts that denote an editing operation, e.g., (a photo of a woman, a photo of a smiling woman), we identify the editing direction in the feature space of diffusion model. Then we align the editing direction of MasterWeaver with that of original T2I model to improve the text controllability without affecting the identity.
@@ 3 Proposed Method
Given one single reference image $x$ of a specific identity, personalized T2I generation aims to generate photo-realistic images of the given identity according to the text prompts $y$. Nonetheless, the generated images are expected to keep a faithful identity and exhibit diversity in attributes, actions, contexts, and so on. To achieve this goal, we propose MasterWeaver, a tuning-free method for generating personalized images with promising identity fidelity and flexible editability. As illustrated in Fig. 2, MasterWeaver first employs an identity mapping network to encode identity feature, and incorporates it with text to guide the image generation. To improve the model editability while keeping identity fidelity, we further propose the id-preserved editability learning, including an editing direction loss and a face-augmented dataset. In the following, we first present an overview of the T2I model utilized in our approach (Sec. 3.1). Then, we introduce the details of the face identity injection (Sec. 3.2) and id-preserved editability learning (Sec. 3.3). Finally, we give our learning objective (Sec. 3.4).
@@ 3.1 Preliminary
In this work, we employ the large-scale pre-trained Stable Diffusion (SD) [68] as our T2I model, which can generate diverse and photo-realistic images. It consists of two components, i.e., the autoencoder $\mathcal{E}(\cdot)$, $\mathcal{D}(\cdot)$) and the conditional diffusion model $\epsilon_{\theta}(\cdot)$. Specifically, the encoder $\mathcal{E}(\cdot)$ is trained to map an image $x$ to a lower dimensional latent space $z=\mathcal{E}(x)$, and the decoder $\mathcal{D}(\cdot)$ is trained to map the latent code back to the image so that $D(\mathcal{E}(x))\approx x$. The conditional diffusion model $\epsilon_{\theta}(\cdot)$ is trained on the latent space of autoencoder, which can generate latent codes based on text condition $y$. The mean-squared loss is employed to train the model:
@@
Cross attention is adopted in SD to steer the generation process by text prompt. Specifically, cross attention adopts the latent image feature $h$ and text feature $\tau(y)$ as input, and transforms them to query $Q=W_{Q}\cdot h$, key $K=W_{K}\cdot\tau(y)$ and value $V=W_{V}\cdot\tau(y)$ by projection layers. $W_{Q}$, $W_{K}$, and $W_{V}$ are weight parameters of query, key, and value projection layers, respectively. Then, attention maps are computed with:
@@ 3.2 Faithful Identity Injection
To leverage the pre-trained SD for generating the personalized images of specified identity, we first introduce a mapping network to encode the identity features, and then integrate the encoded features into the image generation process. As shown in Fig. 2 (b), we employ the pre-trained CLIP image encoder $\psi(\cdot)$ to extract the identity information. To ensure identity fidelity, we utilize the local patch image feature from the penultimate layer of the CLIP model, which contains rich details and is sufficient to represent the target identity faithfully. To bridge the gap between the CLIP features and the SD model, we further employ an identity (ID) mapper $M(\cdot)$, which is trained to project the CLIP features as identity feature $f$:
@@
To effectively integrate the identity feature $f$ into the generation, we adopt the dual cross attention mechanism [90, 85]: $\text{Attention}(Q,K^{f},V^{f})$, where $K^{f}=W^{f}_{K}\cdot f$ and $V^{f}=W^{f}_{V}\cdot f$ represent the projected key and value matrices of the identity information, respectively. $W^{f}_{K}$ and $W^{f}_{V}$ are two additionally introduced projection layers in each cross attention block of SD, which are optimized with our ID mapper simultaneously. To fuse the identity information with the text information, we sum them by:
@@ 3.3 ID-Preserved Editability Learning
Although the proposed method in Sec. 3.2 can generate personalized images with faithful identity, its text controllability is limited. Since the reference identity image and input image used in training are from the same image (as shown in Fig. 4 left), when we train the model under the reconstruction paradigm, the learned identity feature is inevitably entangled with facial attributes (e.g., hair, pose and expression). Such an entanglement weakens the control of the text over generated images, and the model tends to directly copy reference images during generation, resulting in poor editability and diversity. To improve editability while keeping a faithful identity, we propose id-preserved editability learning, which consists of an editing direction loss and a face-augmented dataset.
Editing Direction Loss. To improve the text controllability, we first propose an editing direction loss to facilitate the model learning. The SD model is well-recognized for its superior text controllability and can generate images that closely align with the provided textual descriptions. Intuitively, we can employ its editing capability to regularize MasterWeaver. As shown in Fig. 3, by leveraging paired text prompts that indicate an editing operation, e.g., ($y=$ a photo of a person, $y^{\prime}=$ a photo of a person with <attribute>), we can identify an editing direction in the feature space of diffusion UNet model:
@@
Such an editing direction captures the meaningful prior of SD for semantic editing. Then, we align the editing directions of MasterWeaver with those of SD during training:
@@
where $\epsilon_{\theta^{\prime}}$ denotes our MasterWeaver model and $\text{Sim}(\cdot,\cdot)$ denotes the calculation of cosine similarity. $m$ is the mask of the facial region to ensure the loss is only calculated on the facial area. Obviously, if MasterWeaver neglects the text prompts during generation, its editing directions will be different from those of SD, resulting in a large loss. Besides, the calculation of editing direction is solely based on the text difference, and the encoded semantic information is unrelated to the identity. Therefore, training with our proposed direction loss can improve the model’s editability significantly, with moderate sacrifice of identity fidelity.
Furthermore, our editing direction $\Delta_{\epsilon_{\theta}}(y,y^{\prime})$ can be seen as a variant of DDS [26], which is designed to optimize the edited image by minimizing the delta noise: $\mathcal{L}_{dds}=\|\epsilon_{\theta}(z_{t},t,\tau(y))-\epsilon_{\theta}(\hat{z}_{t},t,\tau(y^{\prime}))\|_{2}^{2}$, where $z$ and $\hat{z}$ are original and edited images, $y$ and $y^{\prime}$ are source and target prompts. In $\mathcal{L}_{edit}$, editing direction $\Delta_{\epsilon_{\theta}}(y,y^{\prime})=\epsilon_{\theta}(z_{t},t,\tau(y^{\prime}))-\epsilon_{\theta}(z_{t},t,\tau(y))$ is used to represent the editing prior of original model. By aligning the editing direction with that of the original model, MasterWeaver can inherit its editability. In our implementation, we use the same $z_{t}$ for both source prompt $y$ and target prompt $y^{\prime}$, and the direction is calculated in feature space instead of noise space. We have collected several attribute-related prompt pairs for training, including hair, age, expression, etc. More details can be found in the Suppl.
Face-Augmented Dataset Construction. As we analyzed, one of the reasons behind the insufficient editability is that the learned model entangles the identity with irrelevant attributes and tends to copy the reference image directly, neglecting the text prompt. To address this issue, we build a face-augmented dataset designed for disentangled identity training. As illustrated in Fig. 4, we employ the E4E [67] and DeltaEdit [53] to perform the attribute editing for the reference identity image. Then, we construct our face-augmented dataset by combining the input image, text prompt, and edited face images. In this dataset, the reference identity image and corresponding input image have the same identity but differ in one specific attribute (e.g., hair color, style, or expression). Such a controlled attribute misalignment can effectively facilitate model learning to extract faithful identity features disentangled with attribute details, thereby improving editability. To maintain identity fidelity, we filter to exclude edited face images with lower face similarity to reference images. In our experiments, we utilize a mix of face-augmented and original datasets for balanced editability and identity fidelity. More details of the dataset are provided in the Suppl.